OpenStudy (thecatman):

3x+5=18 help im demonstrating the site to my teacher

OpenStudy (thecatman):

@jaimelynn67

OpenStudy (thecatman):

dont give the answer, show how to solve

sam (.sam.):

First you gotta isolate the variable 'x' , any ideas?

OpenStudy (skullpatrol):

The addition, subtraction, multiplication, and division properties of equality guarantee that a) adding the same real number to or subtracting the same real number from equal numbers gives equal results b) multiplying or dividing equal numbers by the same nonzero real number gives equal results.

OpenStudy (blowery):

3x + 5 = 18 3x + 5 - 5 = 18 - 5 3x = 13 3x/3 = 13/3 x = 13/3 x = 4 1/3